A farmer wishes to put a diagonal brace on his gate. How long should the brace be if the gate is rectangular 12 by 5 ft
Found 2 solutions by nerdybill, unlockmath:
Answer by nerdybill(7384)

Apply Pythagorean theorem:
Let x = length of diagonal
then
x^2 = 12^2 + 5^2
x^2 = 144 + 25
x^2 = 169
x = 13 feet

Answer by unlockmath(1688)

Hello,
We can use the following formula to solve this. A^2 + B^2 = c^2
A and B will be the sides of the rectangle. Plug in the numbers.
12^2 + 5^2= C^2 This can be rewritten as:
144+ 25 = C^2
169 = C^2 Square root both sides will give us:
C = 13
There we go the length of the diagonal is is 13 feet long.
RJ
Check out a new math book at
www.math.unlock.com